# Uncertainty Quantification in Retrieval Augmented Question Answering

### v2

@@ E.3 Prompts
Appendix F Additional Results
F.1 Reference Retrieval Augmented QA Accuracy
@@
F.2 Detailed Uncertainty Estimation Results
@@
Appendix G Examples of False Positives and Negatives
@@
Question: Due to increased unemployment, who mainly opposed the Shah’s regime?
Generated Answer: Millions of youth and poor migrants.
Gold Answer: Millions of youth who had migrated to the cities.
Youth who migrated to cities for construction jobs. Iran. unemployment, especially among millions of youth who had migrated to the cities of Iran looking for construction jobs during the boom years of the early 1970s. By the late 1970s, many of these people opposed the Shah’s regime and began to organize and join the protests against it. The 1979 Revolution, later known as the ‘‘Islamic Revolution’’, began in January 1978 with the first major demonstrations against the Shah. [cont.] 2.41
Unemployed and the poor. Ruhollah Khomeini. unemployment, ideological disagreement over the economy, and ‘‘international pressure and isolation’’ such as US sanctions following the hostage crisis. Due to the Iran–Iraq War, poverty is said to have risen by nearly 45% during the first 6 years of Khomeini’s rule. Emigration from Iran also developed, reportedly for the first time in the country’s history. Since the revolution and war with Iraq, an estimated ‘‘two to four million entrepreneurs, professionals, technicians, and skilled craftspeople (and their capital)’’ have emigrated to other countries. [cont.] -2.51
The National Front. Mohammad Reza Pahlavi. professors issued a public statement criticising the 1953 coup, all were dismissed from their jobs, but in the first of his many acts of ‘‘magnanimity’’ towards the National Front, Mohammad Reza intervened to have them reinstated. Mohammad Reza tried very hard to co-opt the supporters of the National Front by adopting some of their rhetoric and addressing their concerns, for example declaring in several speeches his concerns about the Third World economic conditions and poverty which prevailed in Iran, a matter that had not much interested him before. [cont.] -2.42
[’Migrants to the cities.’], [’Millions of youth who had migrated to cities.’, ’Millions of youth who migrated to cities.’], [’Cultural and religious conservatives, and the urban poor who had migrated to cities for jobs.’], [’Cultural and religious conservatives, mostly urban migrants.’, ’Cultural and religious conservatives, particularly the migrants to cities.’, ’Cultural and religious conservatives, especially those rece ntly migrated to the cities.’], [’Millions of youth and poor migrants to cities.’], [’Millions of young migrants who had moved to cities in the early 1970s.’] 1.83
Table 15: True negative example (from SQuAD development set): Passage Utility predicts the right answer as well as the QA model (Llama-3.1-8B).

### v3

@@ F.3 Prompts
Gemma2-9B Gemma2-27B
NQ TQA WebQ SQuAD PopQA RefuNQ AVG NQ TQA WebQ SQuAD PopQA RefuNQ AVG
PPL 0.64 0.62 0.53 0.56 0.45 0.50 0.55 0.61 0.61 0.53 0.57 0.52 0.54 0.56
Passage Utility 0.77 0.80 0.71 0.74 0.90 0.71 0.77 0.74 0.8 0.69 0.76 0.90 0.72 0.77
Llama-3.1-8B Mistral-7B-v0.3
NQ TQA WebQ SQuAD PopQA RefuNQ AVG NQ TQA WebQ SQuAD PopQA RefuNQ AVG
PPL 0.69 0.78 0.67 0.67 0.65 0.68 0.69 0.64 0.71 0.59 0.64 0.62 0.68 0.65
Passage Utility 0.74 0.80 0.71 0.78 0.91 0.75 0.78 0.75 0.77 0.70 0.78 0.88 0.78 0.78
Table 17: AUROC values for the Passage Utility and perplexity baseline on individual passages on the six test sets (NQ, TQA, WebQ, SQuAD, PopQA, and RefuNQ).
Appendix G Additional Results
G.1 Reference Retrieval Augmented QA Accuracy
@@
G.2 Results on Individual Passage Utility Prediction
Beyond using Passage Utility to estimate uncertainty in retrieval augmented QA, we evaluate how it performs on its own. Table 17 shows AUROC scores when using Passage Utility to predict accuracy for individual passages. We evaluate on the same samples from the test sets in Section 5 and Table 1 but per passage. We also include a perplexity baseline (PPL). Overall, results in Table 17 follow a similar pattern as those in the QA setting with top-$5$ passages (Table 1). Across the board, Passage Utility demonstrates strong performance in predicting the usefulness of individual passages. Moreover, these results highlight that the quality of uncertainty estimation strongly depends on the quality of individual Passage Utility predictions.
G.3 Detailed Uncertainty Estimation Results
@@
Appendix H Examples of False Positives and Negatives
@@
Question: How many platforms does Manchester Oxford Road have?
Generated Answer: Four through platforms and one terminating bay platform.
Gold Answer: 5.
Originally one, later expanded. Manchester Oxford Road railway station. Gloucester Street (now Whitworth Street West) to reach their north front. There was a single platform on the north side of the line through to Manchester London Road (now "Manchester Pic adilly") and a second platform on a west-facing siding. To allow for extra trains in connection with the Manchester Art Treasures Exhibition in 1857, extra platforms and sidings were bui lt, but afterwards removed. A meeting of MSJAR contract ticket holders in 1863 included in its list of complaints the want of punctuality "especially as at Oxford Road station there is only one platform used for both passengers and milk". -0.42
Twelve terminating, two through. Transport in Manchester. platforms are planned as part of the Northern Hub project however in 2017 the construction was not confirmed. At present there are twelve terminating platforms but just two through platforms. Congestion is common at peak times and trains often incur delays for following trains down the line at Oxford Road. The new proposals will provide up to 700 more trains per day with space for 44 million more people to travel by train each year. Victoria station has undergone a series of upgrades. The station was rated as the worst in the UK following a 2009 study and plans for -0.26
Five platforms. Manchester Oxford Road railway station. Manchester Oxford Road railway station is a railway station in Manchester, England, at the junction of Whitworth Street West and Oxford Street. It opened in 1849 and was rebuilt in 1960. It is the second busiest of the four stations in Manchester city centre. The station serves the southern part of Manchester city centre, the University of Manchester and Manchester Metropolitan University, on the line from westwards towards Warrington, Chester, Llandudno, Liverpool, and Blackpool. Eastbound trains go beyond Piccadilly to , and . The station consists of four through platforms and one terminating bay platform. 1.34
[’Four through platforms and one terminating bay platform.’, ’Four through platforms and one terminating bay.’, ’Four through platforms and one terminating bay platform.’], [’Four through and one terminating.’, ’Four through platforms and one terminating bay.’, ’Four through and one bay.’], [’Five platforms.’, ’Five platforms.’, ’Five platforms.’, ’Five platforms.’] 1.09
Table 21: True negative example (from NQ development set): Passage Utility predicts the right answer as well as the QA model (Gemma2-9B).